OpenStudy (anonymous):

What is the mass of 5.05×1021 platinum atoms?

OpenStudy (abb0t):

First, you need to convert to mols, using avogadros constant: \(\sf \color{blue}{6.022 \times 10^{23}\frac{atoms}{mol}}\), then multiply by the molar mass of Pt to get grams.

OpenStudy (anonymous):

how do I plug my numbers into avogadros constant?

OpenStudy (wolfe8):

1 mole = Avogardo's constant of atoms.

OpenStudy (anonymous):

okay, how can I solve from there?

OpenStudy (wolfe8):

So number of particles in total for platinum(a single atomic element)=nA where n is the number of moles and A is Avogadro's constant.

OpenStudy (anonymous):

So, (5.05 x 10^21) x (6.022 x 10^23) ? Sorry, my lecture class regarding this lesson was cancelled but we still have graded homework due on it. I'm very lost.

OpenStudy (wolfe8):

Translating the equation I gave you: 5.05 x 10^21=n(6.022 x 10^23)

OpenStudy (anonymous):

So, what number plugs into n? The mass number (195.1)?

OpenStudy (wolfe8):

Well you solve for n first to find the number of moles. Here maybe this will help you understand what we are doing http://erhs.rcsnc.org/UserFiles/Servers/Server_4766394/Image/Donna%20Scofield/mole%20conversion%20map.bmp